Sucker hole - winds have limited us to pointing to the north only. High priority ESM target in the north so powering up wavefront sensor

6:30 - wfs_init - housekeeper did not come up

power up ow wavefron sensor - noisy ocam. Cycled BCU2k and cycled binning resolved.

But looks like another wake of clouds will come before we ever get to get on a science target.

7:58 Preset to ESM target, We have some clouds, R=14.25, but estimated magnitude 15.9. Only 1.5 phons/subapp, bin2. We railed in X during close loop. Resending preset with R=15.9. Old PA was 45, rotator 279, new PA 25 actual rotator position 313.

With updated mag of 15.9 - bin4, freq 297, rot pos of 313 we railed in X again.

New PA 82, rotator position 371. Closed. Gains 1.01, 0.78, 0.10. To quote Steve: This program was neither fast, nor easy.

We closed the loop just in time for the clouds to come in.

-- JenniferPower - 09 Mar 2021
Topic revision: r1 - 09 Mar 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback